The area of a rectangle is 24 square centimeters and its side lengths are x centimeters and x + 2 centimenters.
a. find the side lengths of a rectangle. justify your answer
b. find the perimeter of the rectangle

Found 2 solutions by checkley77, advedantatutor:
Answer by checkley77(12844)

AREA=LW
24=X(X+)
24=X^2+2X
X^2+2X-24=0
(X+6)(X-4)=0
X-4=0
X=4 ANS.
PROOF:
24=4(4+2))
24=4*6
24=24

Answer by advedantatutor(4)

Length = x
Width = x + 2
Length x Width = x x (x + 2) = 24
x^2 + 2x = 24
x^2 + 2x - 24 = 0
( x - 4) ( x + 6 ) = 0
x = 4 cm
Length = 4 cm
Width = 4 + 2 = 6 cm
Check :
1) 4 + 2 = 6
2) 4 x 6 = 24
Perimeter = 2 ( Length + width )
= 2 ( 4 + 6 )
= 2 ( 10 )
= 20 cm.